the names, or notes, of all the white keys on a piano are written on pieces of paper and placed in a cup. there are 7 sets of the notes D, E, F, and G and 8 sets of the notes A,B, and C. I f a person draws one note out of the cup without looking, what are his or her chances of getting a c.
Answer by richard1234(7193)

There are 8 pieces of paper with "C" out of 8(3) + 7(4) = 52 total pieces. The probability is 8/52 = 2/13.